OpenStudy (gabylovesyou):

If the cos 25° is 0.906, then the sin _____° is 0.906. Is the answer 50 ?

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (solomonzelman):

sin(A+ {π/2} ) = cos(A) If I am recalling the rule correctly.

OpenStudy (solomonzelman):

Oops I answered the wrong thing

OpenStudy (ipwnbunnies):

It's actually: \[\sin(\frac{π}{2} - x) = \cos(x)\]

OpenStudy (solomonzelman):

:)

OpenStudy (ipwnbunnies):

:)

ganeshie8 (ganeshie8):

also \(\large \frac{\pi}{2} = 90\) so \(\large \sin(90-x) = \cos(x)\)

OpenStudy (gabylovesyou):

thank you

OpenStudy (solomonzelman):

yes same thing

OpenStudy (solomonzelman):

I was talking about sin(90-x) ...